OpenStudy (anonymous):

6x = -7y + 11 5x + 2y = 13

OpenStudy (anonymous):

@Abhisar

OpenStudy (anonymous):

@geekfromthefutur

OpenStudy (anonymous):

@juanpabloJR

OpenStudy (anonymous):

@KlOwNlOvE

OpenStudy (anonymous):

@Preetha

OpenStudy (geekfromthefutur):

What is the question

OpenStudy (anonymous):

6x = -7y + 11 5x + 2y = 13

OpenStudy (geekfromthefutur):

10

OpenStudy (anonymous):

its not a one number answer but thanks for effort

OpenStudy (bradely):

6x = -7y + 11 6x+7y = 11 ... (1) 5x + 2y = 13 ... (2) 5(1)-6(2) 30x+35y=55 - 30x+12y=78 ------------ 23y=-23 y=-1 6x = -7(-1) + 11 6x=18 x=3 source: http://www.mathskey.com/question2answer/